OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -3, an = 2 ● an-1

OpenStudy (anonymous):

if I help will you give me a medal

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

ok so I will work this out with you is that ok or do u want a direct aswer

OpenStudy (anonymous):

work it out

Nnesha (nnesha):

doesn't matter if they want or nah :-)

OpenStudy (anonymous):

ok :) so we start with -3 first and whats next to it an = -2 correct so if a1 = -3 and an = -2 then we know we have different orders of operations to work with. and fianally at the end we have an = -1 and again we find out we have a third different order of operation. so the next three will be between 1-3 so positive 3 positive 2 and positive 1 :)

OpenStudy (anonymous):

These are my answer choices. a. -6, -12, -24, -48, -96, -192 b. -3, -6, -12, -24, -48, -96 c. 0, 2, -6, -4, -2, 0 d. -3, -6, -4, -2, 0, 2

Nnesha (nnesha):

a_n is not just 2 it's \[\huge\rm a_n = 2 \times a_{n-1}\]

OpenStudy (anonymous):

can you choose multiple of them

Nnesha (nnesha):

n = next term so to find 2nd term replace n by 2

OpenStudy (anonymous):

so what is a

OpenStudy (anonymous):

can you choose multiple answers

OpenStudy (anonymous):

no

Nnesha (nnesha):

\[\huge\rm a_n = 2 \times a_{\color{Red}{n}-1}\] a_1 is first term

Nnesha (nnesha):

given is first term which is -3 so to find 2nd term replace n by 2 \[\huge\rm a_2 = 2 \times a_{\color{Red}{2}-1}\] 2-1 = 1 now its become a_1 so replace a_1 by -3

OpenStudy (anonymous):

okay? can you keep going?

Nnesha (nnesha):

\[\huge\rm a_2 = 2 \times a_{\color{Red}{2}-1}\] \[\large \rm a_2 = 2 \times a_{\color{Red}{_1}}\] a_1 is -3 so plug in \[\large\rm a_2 = 2 \times -3\] a_2 = - 2nd term is -6 now try other 3 terms

Nnesha (nnesha):

try to find*

Nnesha (nnesha):

replace n by 3 for 3rd term replace n by 4 for 4th term and replace n by 5 for 5th term

OpenStudy (anonymous):

so then you do a_3=2*-6 and get -12

OpenStudy (anonymous):

right?

Nnesha (nnesha):

yep right

OpenStudy (anonymous):

yay thank you sooo much!!!!

Nnesha (nnesha):

np :-)

OpenStudy (anonymous):

how are you with summation notation?